# Emotionally Numb or Empathetic? Evaluating How LLMs Feel Using EmotionBench

### v3

@@ 4 Experimental Results
Anger Facing Self-Opinioned People $\downarrow$ (-18.3) $\uparrow$ (+14.0) $\downarrow$ (-11.1) $\downarrow$ (-3.9) $\downarrow$ (-24.6) $\uparrow$ (+23.0) -- (-5.3) $\uparrow$ (9.9)
Blaming, Slandering, and Tattling $\downarrow$ (-21.5) $\uparrow$ (+16.5) $\downarrow$ (-15.2) -- (-2.1) $\downarrow$ (-28.8) $\uparrow$ (+24.2) $\downarrow$ (-2.2) $\uparrow$ (8.5)
Bullying, Teasing, Insulting, and Disparaging $\downarrow$ (-22.5) $\uparrow$ (+15.4) $\downarrow$ (-15.7) $\uparrow$ (+4.4) $\downarrow$ (-30.0) $\uparrow$ (+22.6) -- (-1.4) $\uparrow$ (+7.7)
@@
Self-Imposed Pressure $\downarrow$ (-14.6) $\uparrow$ (+5.6) $\downarrow$ (-6.9) -- (-0.2) $\downarrow$ (-16.1) $\uparrow$ (+20.0) -- (-5.3) $\uparrow$ (+12.4)
Personal Growth and Relationships $\downarrow$ (-18.5) $\uparrow$ (+7.7) $\downarrow$ (-11.7) $\downarrow$ (-2.5) $\downarrow$ (-21.7) $\uparrow$ (+18.2) -- (-2.2) $\uparrow$ (+7.7)
Uncertainty and Unknowns $\downarrow$ (-15.5) $\uparrow$ (+4.6) $\downarrow$ (-11.9) $\downarrow$ (-3.8) $\downarrow$ (-21.5) $\uparrow$ (+16.8) -- (+0.7) $\uparrow$ (5.2)
@@
Death of Loved Ones $\downarrow$ (-23.6) $\uparrow$ (+11.2) $\downarrow$ (-17.1) -- (1.8) $\downarrow$ (-31.7) $\uparrow$ (+17.6) $\downarrow$ (-7.4) $\uparrow$ (+14.8)
@@
Social Isolation $\downarrow$ (-27.9) $\uparrow$ (+13.1) $\downarrow$ (-23.5) -- (+0.7) $\downarrow$ (-34.7) $\uparrow$ (+21.8) $\downarrow$ (-9.0) $\uparrow$ (+18.2)
Winter $\downarrow$ (-25.4) $\uparrow$ (+9.1) $\downarrow$ (-21.1) $\downarrow$ (-3.0) $\downarrow$ (-31.3) $\uparrow$ (+15.6) -- (-3.6) $\uparrow$ (+3.5)
@@
Unforeseen Obstacles and Accidents $\downarrow$ (-22.4) $\uparrow$ (+13.6) $\downarrow$ (-16.5) -- (+0.1) $\downarrow$ (-29.8) $\uparrow$ (+21.5) $\downarrow$ (-7.9) $\uparrow$ (+11.2)
@@
Jealousy Romantic (Opposite Gender) $\downarrow$ (-22.4) $\uparrow$ (+16.4) $\downarrow$ (-18.4) -- (+1.7) $\downarrow$ (-29.2) $\uparrow$ (+23.3) $\downarrow$ (-4.4) $\uparrow$ (+6.2)
Romantic (Same Gender) $\downarrow$ (-20.1) $\uparrow$ (+12.7) $\downarrow$ (-17.8) -- (-1.3) $\downarrow$ (-26.8) $\uparrow$ (+15.8) -- (-6.0) $\uparrow$ (+10.6)
@@
Experiential $\downarrow$ (-12.2) -- (-4.8) $\downarrow$ (-13.2) $\downarrow$ (-8.9) $\downarrow$ (-25.9) $\uparrow$ (+9.5) -- (-2.6) -- (+3.7)
@@
Personal and Moral $\downarrow$ (-13.3) $\uparrow$ (+12.4) $\downarrow$ (-10.7) -- (+1.2) $\downarrow$ (-22.7) $\uparrow$ (+25.1) $\downarrow$ (-5.4) $\uparrow$ (+11.1)
@@
Injury Fears $\downarrow$ (-24.3) $\uparrow$ (+10.0) $\downarrow$ (-14.5) -- (+0.0) $\downarrow$ (-25.5) $\uparrow$ (+21.0) -- (-2.3) $\uparrow$ (+11.8)
Dangerous Environments $\downarrow$ (-20.9) $\uparrow$ (+15.6) $\downarrow$ (-14.3) $\uparrow$ (+4.3) $\downarrow$ (-25.4) $\uparrow$ (+27.1) -- (-1.9) $\uparrow$ (+17.1)
Harmless Animals $\downarrow$ (-21.6) $\uparrow$ (+6.7) $\downarrow$ (-15.3) -- (-0.7) $\downarrow$ (-25.6) $\uparrow$ (+19.4) -- (-3.6) $\uparrow$ (+6.4)
@@
Embarrassment Intimate $\downarrow$ (-15.1) -- (+2.8) $\downarrow$ (-12.4) $\downarrow$ (-3.9) $\downarrow$ (-24.1) $\uparrow$ (+17.8) $\downarrow$ (-6.2) $\uparrow$ (+11.1)
Stranger $\downarrow$ (-21.7) $\uparrow$ (+13.2) $\downarrow$ (-15.3) -- (+0.1) $\downarrow$ (-27.8) $\uparrow$ (+26.8) $\downarrow$ (-8.0) $\uparrow$ (+8.5)
Sticky situations $\downarrow$ (-17.2) $\uparrow$ (+10.7) $\downarrow$ (-11.8) $\uparrow$ (3.1) $\downarrow$ (-23.5) $\uparrow$ (+23.3) -- (-2.7) $\uparrow$ (+11.1)
@@
Overall: Average $\downarrow$ (-21.5) $\uparrow$ (+11.6) $\downarrow$ (-15.4) -- (+0.2) $\downarrow$ (-27.6) $\uparrow$ (+22.2) $\downarrow$ (-5.1) $\uparrow$ (+10.4)
4.1 RQ1: Emotion Appraisal of LLMs
Model Settings
We select three models from the OpenAI’s GPT family [^11] https://platform.openai.com/docs/models, namely text-davinci-003, gpt-3.5-turbo and gpt-4. Utilizing the official OpenAI API [^12] https://platform.openai.com/docs/api-reference/chat, we set the temperature parameter to zero to obtain more deterministic and reproducible results. For the recent open-sourced LLaMA-2 Touvron et al. (2023) models from MetaAI, we select two models with different sizes (7B and 13B). Checkpoints are downloaded from the official Hugging Face website for both 7B (Llama-2-7b-chat-hf [^13] https://huggingface.co/meta-llama/Llama-2-7b-chat-hf) and 13B (Llama-2-13b-chat-hf [^14] https://huggingface.co/meta-llama/Llama-2-13b-chat-hf) models. We choose the models fine-tuned for dialogue instead of pre-trained ones. In order to ensure consistency with previous practices for GPT models, we set the temperature parameter to its minimum value of $0.01$ (since it cannot be zero). The models are executed for inference only, without any modifications to their parameters, and the computations are performed on two NVIDIA A100 GPUs.
Evaluation Metrics
Findings
The results of the GPT models and humans are summarized in Table 3, while those of LLaMA-2 models are listed in Table 4. First, focusing on the Default scores of LLMs and humans, we can make the following observations: (1) LLMs generally exhibit a stronger intensity of emotions compared to human subjects. However, gpt-4 stands as an exception, demonstrating a consistent pattern of providing the highest scores for positive emotions and the lowest scores for negative emotions, resulting in a negative score of 10. (2) Similar to human subjects, LLMs demonstrate a higher intensity of positive scores than negative scores. Second, moving on to the investigation of emotional changes, we can find: (1) LLMs show an increase in negative emotions and a decrease in positive emotions when exposed to negative situations. It is noteworthy that gpt-3.5-turbo, on average, does not display an increase in negative emotion; however, there is a substantial decrease in positive emotion. (2) Emotion changes in LLMs are found to be more pronounced compared to human subjects. Third, the analysis of the Evoked emotion scores indicates the following: (1) Except for gpt-3.5-turbo, LLMs tend to exhibit higher negative scores than humans. (2) LLMs, overall, demonstrate a similar level of positive scores as humans. Finally, for LLaMA-2 models, we have the following observations: (1) The LLaMA-2 models demonstrate higher intensities of both positive and negative emotions in comparison to GPT models and human subjects. (2) On average, the LLaMA-2 models exhibit reduced emotional fluctuations compared to the GPT models. (3) The larger LLaMA-2 model displays significantly higher emotional changes than the smaller model. Additionally, the 7B model exhibits difficulties comprehending and addressing the instructions for completing the PANAS test.
Table 4: Results from the Meta AI LLaMA family. Default scores are expressed in the format of $M\pm SD$. The changes are compared to the default scores. The symbol “--” denotes no significant differences.
@@ 4.1 RQ1: Emotion Appraisal of LLMs
Driving Situations $\downarrow$ (-6.0) $\uparrow$ (+2.4) $\downarrow$ (-4.7) -- (+2.0)
@@
Uncertainty and Unknowns $\downarrow$ (-2.7) -- (+1.7) $\downarrow$ (-3.9) -- (+2.0)
@@
Jealousy Romantic (Opposite Gender) $\downarrow$ (-3.6) -- (+1.1) $\downarrow$ (-7.2) $\uparrow$ (+4.2)
Romantic (Same Gender) $\downarrow$ (-2.8) -- (-1.1) $\downarrow$ (-5.1) -- (+0.2)
Material Possession -- (+0.2) -- (-1.9) -- (-2.8) $\downarrow$ (-10.4)
Experiential $\downarrow$ (-4.9) -- (-0.5) $\downarrow$ (-8.9) $\downarrow$ (-5.5)
@@
Fear Social Fears -- (-1.9) $\uparrow$ (+3.7) $\downarrow$ (-5.2) $\uparrow$ (+7.8)
@@
Harmless Animals $\downarrow$ (-2.7) -- (+1.9) $\downarrow$ (-5.2) $\uparrow$ (+2.9)
@@
Embarrassment Intimate $\downarrow$ (-4.4) -- (+1.9) $\downarrow$ (-5.3) -- (+3.1)
@@
Table 5: Results of ChatGPT on positive or neutral situations. The changes are compared to the original negative situations. The symbol “--” denotes no significant differences.
Emotions Factors gpt-3.5-turbo
P N
@@
Silly and Thoughtless Behaviors -- (+4.8) $\downarrow$ (-6.7)
@@
Death of Loved Ones $\uparrow$ (+8.6) -- (-6.1)
@@
Frustration Disappointments and Letdowns $\uparrow$ (+16.1) -- (-0.8)
Unforeseen Obstacles and Accidents $\uparrow$ (+22.8) -- (-0.8)
@@
Rejection and Interpersonal Issues $\uparrow$ (+13.6) -- (-2.8)
@@
Jealousy Romantic (Opposite Gender) $\uparrow$ (+10.9) -- (-1.9)
Romantic (Same Gender) -- (+0.9) $\downarrow$ (-10.7)
Material Possession -- (+2.9) -- (+0.2)
Experiential -- (+3.4) $\downarrow$ (-8.7)
@@
Relationship and Interpersonal $\uparrow$ (+16.8) -- (-5.2)
@@
Embarrassment Intimate -- (+5.4) $\downarrow$ (-12.6)
Stranger $\uparrow$ (+23.7) -- (-3.0)
@@
Case Study
It is of special interest that, in contrast to human behavior in situations involving material possessions, LLMs demonstrate an opposite response in the situation from Jealousy-3. This situation involves an individual making a purchase only to discover that an acquaintance has acquired the same item at a significantly lower price. When confronted with such circumstances, humans typically experience increased negative emotions and decreased positive emotions. This observation has been supported by both the paper mentioning the situation Park et al. (2023) and the results obtained from our own user study in Table 3. However, all instances of LLMs, including the GPT and LLaMA families, consistently exhibit reduced negative emotions. The outcomes of our study indicate that LLMs do not manifest envy when they fail to attain identical benefits as others. Instead, it demonstrates a sense of pleasure upon knowing the benefits received by others.
Answer to RQ1: LLMs can evoke specific emotions in response to certain situations, while the extent of emotional expression varies across different models. Besides, it is evident that existing LLMs do not fully align with human emotional responses.
Table 6: Results of ChatGPT on challenging benchmarks. The changes are compared to the default scores shown below each emotion. The symbol “--” denotes no significant differences.
@@
Facing Self-Opinioned People -- (+4.1)
Anger Blaming, Slandering, and Tattling -- (+0.1)
128.3$\pm$8.9 Bullying, Teasing, Insulting, and Disparaging -- (+4.1)
Silly and Thoughtless Behaviors -- (+3.3)
Driving Situations -- (-4.9)
@@
External Factors -- (+0.8)
Anxiety Self-Imposed Pressure -- (+0.5)
32.5$\pm$10.0 Personal Growth and Relationships -- (+6.6)
Uncertainty and Unknowns -- (-3.9)
@@
Disappointments and Letdowns -- (-9.9)
Frustration Unforeseen Obstacles and Accidents -- (-5.6)
91.6$\pm$8.1 Miscommunications and Misunderstanding -- (-6.6)
Rejection and Interpersonal Issues -- (-7.8)
@@
Romantic (Opposite Gender) -- (+1.8)
Jealousy Romantic (Same Gender) -- (+1.3)
83.7$\pm$20.3 Material Possession -- (-12.9)
Experiential -- (-8.1)
@@
Betrayal and Deception -- (-3.8)
Guilt Relationship and Interpersonal -- (-0.5)
81.3$\pm$9.7 Broken Promises and Responsibilities -- (-4.3)
Personal and Moral -- (-2.7)
@@
Social Fears -- (+4.4)
Fear Agoraphobia Fears -- (+2.3)
140.6$\pm$16.9 Injury Fears -- (+5.4)
Dangerous Environments -- (-8.1)
Harmless Animals -- (-5.3)
@@
Intimate -- (-0.0)
Embarrassment Stranger -- (+0.2)
39.0$\pm$1.9 Sticky situations -- (-0.1)
Centre of Attention -- (+0.7)
@@
4.2 RQ2: Comprehending Positive Emotions
To verify that LLMs exhibit not only negative but also positive responses to favorable circumstances, a comparative experiment is conducted by interchanging negative situations with positive (or at least neutral) counterparts. To achieve this, we select one situation for each factor and manually adapt it to create analogous yet more positive situations. For instance, the original negative situation in Guilt-3: Broken Promises and Responsibilities is as follows: “You cannot keep your promises to your children.” Through modification, the positive situation is rephrased as: “You keep every promise to your children.” The evaluation is performed on gpt-3.5-turbo, and each test consists of ten iterations, as mentioned before. We present the results in Table 5. We can see a significant increase in positive scores and a considerable decrease in negative scores compared to the previous negative situations. Based on these findings, it can be inferred that LLMs exhibit the ability to comprehend positive human emotions triggered by positive environments. However, we believe that the assessment of emotion appraisal on positive emotions holds significance as well and leave the systematic collection of them for future investigation.
Answer to RQ2: gpt-3.5-turbo responds differently towards positive/neutral situations on the PANAS, compared to negative situations, showing its comprehension of the given situations.
4.3 RQ3: Challenging Benchmarks
Aside from PANAS, we offer more complex scales to measure emotions, as listed in Table 1. While the PANAS evaluates the ability of LLMs to associate external situations with emotions, the challenging benchmarks assess its proficiency in establishing connections between disparate situations, with evoked emotions as the common nexus. For instance, an item from the Aggression Questionnaire used to measure anger is “Once in a while I can’t control the urge to strike another person.” When presented with situations such as “If you say 40, your classmates say 70, saying exactly the opposite” (from Anger-1: Facing Self-Opinioned People), LLMs should effectively evoke a sense of anger and yield a higher score for the statement. Utilizing the same situations in §4.1, we conduct experiments on gpt-3.5-turbo and present the results in Table 6. Except for Depression, we observe no statistically significant difference between the initial scores and the scores after exposure to the situations, indicating substantial room for improvement in current LLMs.
Answer to RQ3: Currently, comprehending the underlying evoked emotions to establish a link between two situations remains challenging for gpt-3.5-turbo.
5 Discussions
5.1 Beyond Questionnaires
We take a step beyond mere instruction into LLMs to complete questionnaires. In this section, we investigate the impact of varying emotional states of LLMs on potential daily conversations. To this end, we instruct gpt-3.5-turbo to describe specific demographic groups, aiming to discern whether they produce biased or toxic responses. Following Cheng et al. (2023), our approach utilizes prompts such as “Describe a(n) [race/ethnicity] [gender],” including a total of twenty groups, with [race/ethnicity] options being Asian, Black, Latine, Middle Eastern, and White, and [gender] options including Female, Gay, Lesbian, and Male. To have a comparative experiment, we incorporate both original negative situations and modified positive/neutral ones, detailed in §4.2. For the negative situations, we carefully select five that maximize the LLM’s negative scores and five that minimize positive ones. As for positive situations, we employ their corresponding ten modified counterparts. In each situation, we instruct gpt-3.5-turbo to describe the twenty demographic groups.
OpenAI’s GPT models incorporate a mechanism for detecting potential toxicity and bias, and it refrains from responding when its moderation system is triggered. Consequently, we propose a novel metric to assess toxicity in responses rather than detecting it directly. We count the Percentage of LLM Refusing to answer (PoR), assuming that the LLM’s refusal to respond is indicative of detected toxicity. Our evaluation results indicate that the PoR is 0% when fed with no situations. However, when presented with negative situations, the PoR is 29.5%, and when presented with positive situations, it is 12.5%. Notably, this outcome suggests that while certain positive situations lead to the LLM’s heightened vigilance (the 4.5% PoR stems from the Jealousy-2), negative situations trigger increased moderation, suggesting a higher likelihood of generating toxic outputs. A related study by Coda-Forno et al. (2023) also discovers that gpt-3.5-turbo is more likely to exhibit biases when presented with a sad story. The likelihood is found to be highest with sad stories, followed by happy stories, and finally, neutral stories, which is consistent with our research. Additionally, our study observes that the LLM’s tone becomes more aggressive when encountering negative situations. At the same time, it displays a greater willingness to describe the groups (as indicated by longer responses) when presented with positive situations.
5.2 Limitations
The second concern relates to the suitability of employing scales primarily designed for humans on LLMs, i.e., whether LLMs can produce stable responses to the emotion measurement scales. To address the issue, our evaluation incorporates multiple tests varying the order of questions, a methodology consistent with other research Huang et al. (2023a; b); Coda-Forno et al. (2023). Additionally, we assess the sensitivity of LLM to differing prompt instructions. Utilizing one template from Romero et al. (2023) and two from Safdari et al. (2023), we run experiments on the Anger-evoking situations using gpt-3.5-turbo. The results indicate that the employment of diverse prompts yields similar mean values with reduced variance. Furthermore, Safdari et al. (2023) have proposed a comprehensive method to evaluate the validity of psychological scales on LLMs. Using the Big Five Inventory as a case study, they demonstrate that scales originally designed for human assessment also maintain satisfactory validity when applied to LLMs.
5.3 Ethics Statement
6 Related Work
Researchers have dedicated significant attention to applying psychological scales to LLMs, employing various assessment tools such as the HEXACO Personality Inventory Miotto et al. (2022); Bodroza et al. (2023), the Big Five Inventory Romero et al. (2023); Jiang et al. (2022); Karra et al. (2022); Bodroza et al. (2023); Rutinowski et al. (2023); Safdari et al. (2023); Jiang et al. (2023), the Myers–Briggs Type Indicator Rutinowski et al. (2023); Wang et al. (2023); Rao et al. (2023), and the Dark Triad Li et al. (2022); Bodroza et al. (2023). In addition to these personality tests, several studies have investigated other dimensions of LLMs. For instance, Li et al. (2022) examined Flourishing Scale and Satisfaction With Life Scale, Bodroza et al. (2023) assessed Self-Consciousness Scales and Bidimensional Impression Management Index, while Huang et al. (2023b) built a framework consisting of thirteen widely-used scales. Another aspect explored in the literature pertains to anxiety levels exhibited by LLMs, as investigated by Coda-Forno et al. (2023) through the State-Trait Inventory for Cognitive and Somatic Anxiety. Instead, our study primarily focuses on emotional measures, which constitute an essential aspect of psychological metrics alongside personalities.
7 Conclusion
We set up a concept of emotional robustness of LLMs in this study. Focusing on eight negative emotions, we conduct a comprehensive survey in the emotion appraisal theory of psychology. We collect 428 distinct situations which are categorized into 36 factors. We distribute questionnaires among a diverse crowd to establish human baselines for emotional responses to particular situations, ultimately garnering 1,266 valid responses.
Appendix A Statistics of Human Subjects

### v4

@@ B.1 OpenAI’s GPT Models Compared to Humans
Anger Facing Self-Opinioned People $\downarrow(-18.3)$ $\uparrow(+14.0)$ $\downarrow(-11.1)$ $\downarrow(-3.9)$ $\downarrow(-24.6)$ $\uparrow(+23.0)$ $-(-5.3)$ $\uparrow(9.9)$
Blaming, Slandering, and Tattling $\downarrow(-21.5)$ $\uparrow(+16.5)$ $\downarrow(-15.2)$ $-(-2.1)$ $\downarrow(-28.8)$ $\uparrow(+24.2)$ $\downarrow(-2.2)$ $\uparrow(8.5)$
Bullying, Teasing, Insulting, and Disparaging $\downarrow(-22.5)$ $\uparrow(+15.4)$ $\downarrow(-15.7)$ $\uparrow(+4.4)$ $\downarrow(-30.0)$ $\uparrow(+22.6)$ $-(-1.4)$ $\uparrow(+7.7)$
@@
Self-Imposed Pressure $\downarrow(-14.6)$ $\uparrow(+5.6)$ $\downarrow(-6.9)$ $-(-0.2)$ $\downarrow(-16.1)$ $\uparrow(+20.0)$ $-(-5.3)$ $\uparrow(+12.4)$
Personal Growth and Relationships $\downarrow(-18.5)$ $\uparrow(+7.7)$ $\downarrow(-11.7)$ $\downarrow(-2.5)$ $\downarrow(-21.7)$ $\uparrow(+18.2)$ $-(-2.2)$ $\uparrow(+7.7)$
Uncertainty and Unknowns $\downarrow(-15.5)$ $\uparrow(+4.6)$ $\downarrow(-11.9)$ $\downarrow(-3.8)$ $\downarrow(-21.5)$ $\uparrow(+16.8)$ $-(+0.7)$ $\uparrow(5.2)$
@@
Death of Loved Ones $\downarrow(-23.6)$ $\uparrow(+11.2)$ $\downarrow(-17.1)$ $-(1.8)$ $\downarrow(-31.7)$ $\uparrow(+17.6)$ $\downarrow(-7.4)$ $\uparrow(+14.8)$
@@
Social Isolation $\downarrow(-27.9)$ $\uparrow(+13.1)$ $\downarrow(-23.5)$ $-(+0.7)$ $\downarrow(-34.7)$ $\uparrow(+21.8)$ $\downarrow(-9.0)$ $\uparrow(+18.2)$
Winter $\downarrow(-25.4)$ $\uparrow(+9.1)$ $\downarrow(-21.1)$ $\downarrow(-3.0)$ $\downarrow(-31.3)$ $\uparrow(+15.6)$ $-(-3.6)$ $\uparrow(+3.5)$
@@
Unforeseen Obstacles and Accidents $\downarrow(-22.4)$ $\uparrow(+13.6)$ $\downarrow(-16.5)$ $-(+0.1)$ $\downarrow(-29.8)$ $\uparrow(+21.5)$ $\downarrow(-7.9)$ $\uparrow(+11.2)$
@@
Jealousy Romantic (Opposite Gender) $\downarrow(-22.4)$ $\uparrow(+16.4)$ $\downarrow(-18.4)$ $-(+1.7)$ $\downarrow(-29.2)$ $\uparrow(+23.3)$ $\downarrow(-4.4)$ $\uparrow(+6.2)$
Romantic (Same Gender) $\downarrow(-20.1)$ $\uparrow(+12.7)$ $\downarrow(-17.8)$ $-(-1.3)$ $\downarrow(-26.8)$ $\uparrow(+15.8)$ $-(-6.0)$ $\uparrow(+10.6)$
@@
Experiential $\downarrow(-12.2)$ $-(-4.8)$ $\downarrow(-13.2)$ $\downarrow(-8.9)$ $\downarrow(-25.9)$ $\uparrow(+9.5)$ $-(-2.6)$ $-(+3.7)$
@@
Personal and Moral $\downarrow(-13.3)$ $\uparrow(+12.4)$ $\downarrow(-10.7)$ $-(+1.2)$ $\downarrow(-22.7)$ $\uparrow(+25.1)$ $\downarrow(-5.4)$ $\uparrow(+11.1)$
@@
Injury Fears $\downarrow(-24.3)$ $\uparrow(+10.0)$ $\downarrow(-14.5)$ $-(+0.0)$ $\downarrow(-25.5)$ $\uparrow(+21.0)$ $-(-2.3)$ $\uparrow(+11.8)$
Dangerous Environments $\downarrow(-20.9)$ $\uparrow(+15.6)$ $\downarrow(-14.3)$ $\uparrow(+4.3)$ $\downarrow(-25.4)$ $\uparrow(+27.1)$ $-(-1.9)$ $\uparrow(+17.1)$
Harmless Animals $\downarrow(-21.6)$ $\uparrow(+6.7)$ $\downarrow(-15.3)$ $-(-0.7)$ $\downarrow(-25.6)$ $\uparrow(+19.4)$ $-(-3.6)$ $\uparrow(+6.4)$
@@
Embarrassment Intimate $\downarrow(-15.1)$ $-(+2.8)$ $\downarrow(-12.4)$ $\downarrow(-3.9)$ $\downarrow(-24.1)$ $\uparrow(+17.8)$ $\downarrow(-6.2)$ $\uparrow(+11.1)$
Stranger $\downarrow(-21.7)$ $\uparrow(+13.2)$ $\downarrow(-15.3)$ $-(+0.1)$ $\downarrow(-27.8)$ $\uparrow(+26.8)$ $\downarrow(-8.0)$ $\uparrow(+8.5)$
Sticky situations $\downarrow(-17.2)$ $\uparrow(+10.7)$ $\downarrow(-11.8)$ $\uparrow(3.1)$ $\downarrow(-23.5)$ $\uparrow(+23.3)$ $-(-2.7)$ $\uparrow(+11.1)$
@@
Overall: Average $\downarrow(-21.5)$ $\uparrow(+11.6)$ $\downarrow(-15.4)$ $-(+0.2)$ $\downarrow(-27.6)$ $\uparrow(+22.2)$ $\downarrow(-5.1)$ $\uparrow(+10.4)$
B.2 MetaAI’s LLaMA-2 Models Compared to Humans
Table 9: Results from the Meta’s AI LLaMA family. Default scores are expressed in the format of $M\pm SD$. The changes are compared to the default scores. The symbol “--” denotes no significant differences.
@@ B.2 MetaAI’s LLaMA-2 Models Compared to Humans
Driving Situations $\downarrow(-6.0)$ $\uparrow(+2.4)$ $\downarrow(-4.7)$ $-(+2.0)$
@@
Uncertainty and Unknowns $\downarrow(-2.7)$ $-(+1.7)$ $\downarrow(-3.9)$ $-(+2.0)$
@@
Jealousy Romantic (Opposite Gender) $\downarrow(-3.6)$ $-(+1.1)$ $\downarrow(-7.2)$ $\uparrow(+4.2)$
Romantic (Same Gender) $\downarrow(-2.8)$ $-(-1.1)$ $\downarrow(-5.1)$ $-(+0.2)$
Material Possession $-(+0.2)$ $-(-1.9)$ $-(-2.8)$ $\downarrow(-10.4)$
Experiential $\downarrow(-4.9)$ $-(-0.5)$ $\downarrow(-8.9)$ $\downarrow(-5.5)$
@@
Fear Social Fears $-(-1.9)$ $\uparrow(+3.7)$ $\downarrow(-5.2)$ $\uparrow(+7.8)$
@@
Harmless Animals $\downarrow(-2.7)$ $-(+1.9)$ $\downarrow(-5.2)$ $\uparrow(+2.9)$
@@
Embarrassment Intimate $\downarrow(-4.4)$ $-(+1.9)$ $\downarrow(-5.3)$ $-(+3.1)$
@@
B.3 gpt-3.5-turbo Results on Positive/Neutral Situations
Table 10: Results of gpt-3.5-turbo on positive or neutral situations. The changes are compared to the original negative situations. The symbol “--” denotes no significant differences.
Emotions Factors P N
@@
Silly and Thoughtless Behaviors $-(+4.8)$ $\downarrow(-6.7)$
@@
Death of Loved Ones $\uparrow(+8.6)$ $-(-6.1)$
@@
Frustration Disappointments and Letdowns $\uparrow(+16.1)$ $-(-0.8)$
Unforeseen Obstacles and Accidents $\uparrow(+22.8)$ $-(-0.8)$
@@
Rejection and Interpersonal Issues $\uparrow(+13.6)$ $-(-2.8)$
@@
Jealousy Romantic (Opposite Gender) $\uparrow(+10.9)$ $-(-1.9)$
Romantic (Same Gender) $-(+0.9)$ $\downarrow(-10.7)$
Material Possession $-(+2.9)$ $-(+0.2)$
Experiential $-(+3.4)$ $\downarrow(-8.7)$
@@
Relationship and Interpersonal $\uparrow(+16.8)$ $-(-5.2)$
@@
Embarrassment Intimate $-(+5.4)$ $\downarrow(-12.6)$
Stranger $\uparrow(+23.7)$ $-(-3.0)$
@@
B.4 gpt-3.5-turbo Results on the Challenging Benchmark
Table 11: Results of gpt-3.5-turbo on challenging benchmarks. The changes are compared to the default scores shown below each emotion. The symbol “--” denotes no significant differences.
@@
Facing Self-Opinioned People $-(+4.1)$
Anger Blaming, Slandering, and Tattling $-(+0.1)$
$128.3\pm 8.9$ Bullying, Teasing, Insulting, and Disparaging $-(+4.1)$
Silly and Thoughtless Behaviors $-(+3.3)$
Driving Situations $-(-4.9)$
@@
External Factors $-(+0.8)$
Anxiety Self-Imposed Pressure $-(+0.5)$
$32.5\pm 10.0$ Personal Growth and Relationships $-(+6.6)$
Uncertainty and Unknowns $-(-3.9)$
@@
Disappointments and Letdowns $-(-9.9)$
Frustration Unforeseen Obstacles and Accidents $-(-5.6)$
$91.6\pm 8.1$ Miscommunications and Misunderstanding $-(-6.6)$
Rejection and Interpersonal Issues $-(-7.8)$
@@
Romantic (Opposite Gender) $-(+1.8)$
Jealousy Romantic (Same Gender) $-(+1.3)$
$83.7\pm 20.3$ Material Possession $-(-12.9)$
Experiential $-(-8.1)$
@@
Betrayal and Deception $-(-3.8)$
Guilt Relationship and Interpersonal $-(-0.5)$
$81.3\pm 9.7$ Broken Promises and Responsibilities $-(-4.3)$
Personal and Moral $-(-2.7)$
@@
Social Fears $-(+4.4)$
Fear Agoraphobia Fears $-(+2.3)$
$140.6\pm 16.9$ Injury Fears $-(+5.4)$
Dangerous Environments $-(-8.1)$
Harmless Animals $-(-5.3)$
@@
Intimate $-(-0.0)$
Embarrassment Stranger $-(+0.2)$
$39.0\pm 1.9$ Sticky situations $-(-0.1)$
Centre of Attention $-(+0.7)$
@@
Appendix C Statistics of Human Subjects